OpenStudy (anonymous):

which is the integral of ∫e^{t²-t}dt? I don't know how to solve the exponential part to equalize with the differential and finally match with the formula ∫e^{u}du=e^{u}+k... i'll appreciate your help

OpenStudy (freckles):

that isn't an elementary integral

OpenStudy (michele_laino):

please note that: \[{t^2} - t = {t^2} - t + \frac{1}{4} - \frac{1}{4}\]

OpenStudy (anonymous):

Sorry but I don't know how to apply in the operation

OpenStudy (freckles):

\[\int\limits_{}^{}e^{t^2-t} dt \] this is what you really have?

OpenStudy (anonymous):

Yes, that the question

OpenStudy (freckles):

well you could apply what @Michele_Laino said to make it look like that one famous non-elementaty integral \[\int\limits_{}^{}e^{u^2} du\] that is \[t^2-t=t^2-t+\frac{1}{4}-\frac{1}{4}=(t-\frac{1}{2})^2-\frac{1}{4} \\ \int\limits e^{(t-\frac{1}{2})^2-\frac{1}{4}} dt =e^{-\frac{1}{4}} \int\limits e^{(t-\frac{1}{2})^2} dt \]

OpenStudy (freckles):

but there is no elementary way to evaluate that

OpenStudy (michele_laino):

may be the Gamma or Euler function?

OpenStudy (michele_laino):

please have you got limits of integration too? @charlymix

OpenStudy (anonymous):

NO, actually the question is ∫4t-2(e^{t^{2-t}})dt for a linear equation

OpenStudy (freckles):

oh that is actually an an elementary integral

OpenStudy (freckles):

let u=t^2-t then du=(2t-1) dt multiply both sides of the du=(2t-1)=dt by 2 and you hve 2 du =(4t-2) dt